OpenStudy (anonymous):

An urn contains three red balls, five white balls, and two black balls. Three balls are drawn from the urn at random without replacement. For each red ball drawn, you win $10, and for each black ball drawn, you lose $15. Let X represent your net winnings. Compute E(X), your expected net winnings. E(X) = ?

OpenStudy (anonymous):

\[\mathbb{E}(X)=\sum_{\text{all balls}}(\text{probability of getting certain color})(\text{earnings/losses})\]

OpenStudy (anonymous):

So what's the probability of drawing a red ball? a white ball? a black ball? How much money do you earn/lose for each color?

OpenStudy (anonymous):

red ball probability is 3 out of 10 earn 10$ white ball of 5 out of 10 0 black ball is 2 out of 10 loose 15

OpenStudy (anonymous):

E(x)= 10(3/10) + 0(5/10) + -15(2/10) = 3 + 0 + -3 = 0 is this right?

OpenStudy (anonymous):

Yes